OpenStudy (anonymous):

scott has 4 times as many green marbles as blue marbles. he has a total of 20 marbles. how many more green marbles does scott have the blue marbles?

OpenStudy (anonymous):

no the anwer has to be 5,10,12,or 16

OpenStudy (anonymous):

g = green b = blue g + b = 20 g = 4b plug the second row in first row 4b + b = 20 --> b = 4 plug b in second row g = 4(4) = 16 answer is g - b = 12 more green than blue

OpenStudy (anonymous):

thank u i understand now

OpenStudy (anonymous):

a hobby shop has 4 times as many model airplanes as model helicopters. the total number of models is 35. how many model airplanes does the hobby shop have?

OpenStudy (anonymous):

same thing a + h = 35 a = 4h

OpenStudy (anonymous):

how many model airplanes does the hobby shop have?

OpenStudy (anonymous):

curtis has 36 pennies to put in 4 piles. he writes 36 is times as many as 9. what other comparison sentence can he write to show the comparision